# Rotatable Antenna Aided Mixed Near-Field and Far-Field Communications in the Upper Mid-Band: Interference Analysis and Joint Optimization

### v2

@@ II-A2 Near-field channel model
For a typical near-field user, the channel is modeled based on spherical-wave propagation. Consider the near-field user $k\in\mathcal{K}$, located at a distance of $r_{k}$ and a spatial angle of $\theta_{k}$ from the center of the entire ULA. The distance from the $\bar{n}$-th antenna in subarray $q$ to user $k$, accounting for the subarray’s rotation, is:
$r^{(\bar{n})}_{k}$ $=\sqrt{r_{k}^{2}+\delta^{2}_{q,\bar{n}}-2r_{k}\delta_{q,\bar{n}}(\cos{\phi_{q}}\cos{\theta_{k}}+\sin{\phi_{q}}\sin{\theta_{k}})}$
$\stackrel{{\scriptstyle(a)}}{{\approx}}r_{k}-\delta_{q,\bar{n}}\cos{(\phi_{q}-\theta_{k})}+\frac{\delta^{2}_{q,\bar{n}}(1-\cos^{2}{(\phi_{q}-\theta_{k})})}{2r_{k}},$ (1)
where $(a)$ is due to the second-order Taylor approximation $\sqrt{1+x}\approx 1+\frac{1}{2}x-\frac{1}{8}x^{2}$. Then, the near-field channel steering vector $\mathbf{b}_{q}(\theta_{k},r_{k},\phi_{q})\in\mathbb{C}^{\bar{N}}$ for subarray $q$ is given by
$\mathbf{b}_{q}(\theta_{k},r_{k},\phi_{q})=\left[e^{-j\frac{2\pi}{\lambda}(r^{(\bar{n})}_{k}-r_{k})}\right]_{n\in\mathcal{\bar{N}},q\in\mathcal{Q}}.$ (2)
@@
where $\bm{\phi}=[\phi_{-\tilde{Q}},\cdots,\phi_{\tilde{Q}}]^{T}$ denotes the vector of rotation angles corresponding to all subarrays. In this paper, we adopt a multipath channel model for near-field users, comprising a line-of-sight (LoS) path and $L_{{\rm N},k}$ non-line-of-sight (NLoS) paths induced by environment scatterers. Then, the channel between the BS and the near-field user $k$ is modeled as
$\mathbf{h}^{H}_{{\rm N},k}=\sqrt{N}{h}_{{\rm N},k}\mathbf{b}^{H}(\theta_{k},r_{k},\bm{\phi})+\sum^{L_{{\rm N},k}}_{\ell=1}\sqrt{N}{h}_{{\rm N},\ell}\mathbf{b}^{H}(\theta_{\ell},r_{\ell},\bm{\phi}),$ (4)
where ${h}_{{\rm N}}$ and ${h}_{{\rm N},\ell}$ denote the complex-valued channel gains of the LoS path and the $\ell$-th NLoS path between the BS and near-field user $k$, respectively. [^3] For simplicity, we assume an omni-directional antenna gain model. Nonetheless, the proposed scheme can be readily extended to the general directional antenna model described in [26, 29]. $\theta_{\ell}$ and $r_{\ell}$ are the angle and distance of the $\ell$-th scatterer relative to the center of the ULA.
@@ II-A3 Far-field channel model
The channel between the BS and the far-field user $m\in\mathcal{M}$ follows the planar-wave propagation assumption. Similarly, the far-field channel steering vector $\mathbf{a}_{q}(\psi_{m},\phi_{q})\in\mathbb{C}^{\bar{N}}$ for subarray $q$ is given by
$\mathbf{a}_{q}(\psi_{m},\phi_{q})=\left[e^{-j\frac{2\pi}{\lambda}\delta_{q,\bar{n}}\cos{(\psi_{m}-\phi_{q})}}\right]_{\bar{n}\in\mathcal{\bar{N}},q\in\mathcal{Q}},$ (5)
where $\psi_{m}$ denotes the spatial angle of the far-field user $m$. Accordingly, the overall far-field channel steering vector is given by
@@
Similarly, the multipath far-field channel between the BS and the far-field user $m\in\mathcal{M}$ can be expressed as
$\mathbf{h}^{H}_{{\rm F},m}=\sqrt{N}{h}_{{\rm F},m}\mathbf{a}^{H}(\psi_{m},\phi)+\sum^{L_{{\rm F},m}}_{\ell=1}\sqrt{N}{h}_{{\rm F},\ell}\mathbf{a}^{H}(\psi_{\ell},\phi),$ (7)
where ${h}_{{\rm F},m}$ and ${h}_{{\rm F},\ell}$ denote the complex-valued channel gains of the LoS path and the $\ell$-th NLoS path between the BS and user $m$, respectively.
@@ II-B Signal Model
Let $s_{{\rm N},k}\in\mathbb{C}$ and $s_{{\rm F},m}\in\mathbb{C}$ represent the signals transmitted by the BS to the $k$-th near-field user with power $P_{{\rm N},k}$ and to the $m$-th far-field user with power $P_{{\rm F},m}$, respectively. For each legacy far-field user, the preconfigured beamforming vector is defined as $\mathbf{w}_{{\rm F},m}\in\mathbb{C}^{N},m\in\mathcal{M}$. For the $K$ near-field users, to reduce power consumption and hardware cost in upper mid-band MIMO systems, a hybrid beamforming architecture is employed to serve $K$ near-field users with $K$ $(K\ll N)$ radio frequency (RF) chains [13]. The transmitted signal vector is thus given by $\bar{\mathbf{s}}=\mathbf{F}_{\rm A}\mathbf{F}_{\rm D}\mathbf{s}$, where $\bar{\mathbf{s}}=[s_{{\rm N},1},\ldots,s_{{\rm N},K}]^{T}$, $\mathbf{F}_{\rm D}\in\mathbb{C}^{K\times K}$ is the digital precoder, and $\mathbf{F}_{\rm A}=[\mathbf{v}_{{\rm N},1},\ldots,\mathbf{v}_{{\rm N},K}]\in\mathbb{C}^{N\times K}$ is the analog precoder, with $\mathbf{v}_{{\rm N},k}\in\mathbb{C}^{N}$ representing the analog beamforming vector for the $k$-th near-field user. As such, the received signal at the $k$-th near-field user is given by [^4] To fully exploit the performance gains brought by RAs, we assume perfect channel state information (CSI) for both near-field and far-field users is available at the BS [26]. In practice, their CSI can be effectively acquired through existing far-field and near-field beam training [31] and/or channel estimation techniques [32].
@@
where $\mathbf{w}_{{\rm N},i}=\mathbf{F}_{\rm A}\mathbf{f}_{{\rm D},i}$ with $\mathbf{f}_{{\rm D},i}$ being the $i$-th column of $\mathbf{F}_{\rm D}$. $n_{k}\sim\mathcal{CN}(0,\sigma^{2}_{k})$ represents the additive white Gaussian noise (AWGN) at the near-field user $k$. Accordingly, the achievable rate (bits/s/Hz) of user $k$ is given by $R_{k}=\log_{2}\left(1+\gamma_{k}\right)$, where
@@ II-C Problem Formulation
To shed important insights into the impact of RAs on mixed-field communication systems, we employ a low-complexity hybrid beamforming architecture [13]. In this architecture, $K$ beams are directed to the near-field users using the near-field beam-focusing property, while multiple preconfigured maximum ratio transmission (MRT)-like beams are steered toward the legacy far-field users [12]. The beamforming vectors are defined as $\mathbf{v}_{{\rm N},k}=\frac{\mathbf{h}_{{\rm N},k}}{\|\mathbf{h}_{{\rm N},k}\|}$ for the $k$-th near-field user and $\mathbf{w}_{{\rm F},m}=\frac{\mathbf{h}_{{\rm F},m}}{\|\mathbf{h}_{{\rm F},m}\|}$ for the $m$-th far-field user. The digital beamforming is configured as an identity matrix, i.e., $\mathbf{F}_{\rm D}=\mathbf{I}_{K}$ [13, 33]. [^5] To enhance the system performance, the weighted minimum mean square error (WMMSE) or zero-forcing (ZF) based digital beamforming can be properly designed using the analog precoder $\mathbf{F}_{\rm A}$ [13, 33], with performance evaluations presented in Section V. Consequently, the achievable rate of the near-field user $k$ is given in (II-B) at the bottom of the next page. Within this framework, we consider a system sum-rate maximization problem by jointly optimizing the power allocation $\{P_{{\rm N},k}\}$ and subarray rotation angles $\bm{\phi}$. The optimization problem is formulated as follows:
$({\bf P1}):\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \max_{\begin{subarray}{c}\{P_{{\rm N},k}\},\bm{\phi}\end{subarray}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{k\in\mathcal{K}}\leavevmode\nobreak\ R_{k}\left(\{P_{{\rm N},k}\},\bm{\phi}\right)$ (11a)
s.t. $\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{k\in\mathcal{K}}P_{{\rm N},k}\leq P,$ (11b)
$\leavevmode\nobreak\ \leavevmode\nobreak\ \left|[\bm{\phi}]_{p}-[\bm{\phi}]_{q}\right|>\frac{\pi}{2},\leavevmode\nobreak\ \forall{p,q\in\mathcal{Q}},$ (11c)
$\leavevmode\nobreak\ \leavevmode\nobreak\ [\bm{\phi}]_{q}\in\mathcal{C}_{{\phi_{q}}},\leavevmode\nobreak\ \forall{q\in\mathcal{Q}},$ (11d)
@@
where (11b) enforces the BS transmit power constraint, with $P$ being the maximum transmit power. (11c) ensures that any two subarrays form an obtuse angle after their individual rotations to avoid mutual signal reflections [21, 26]. (11d) restricts the rotation angle of subarray $q$ to its corresponding admissible range $\mathcal{C}_{\phi_{q}}=[\phi_{q,{\rm min}},\phi_{q,{\rm max}}]$.
@@
The optimization problem (P1) is non-convex due to the following factors: 1) the objective function is not jointly concave in terms of the optimization variables $\{P_{{\rm N},k}\}$ and $\bm{\phi}$; 2) these variables are intricately coupled within the objective function; and 3) the rotation angles $\bm{\phi}$ introduce complex dependencies in both near-field and far-field steering vectors. These characteristics render the optimal solution of (P1) computationally challenging. To tackle this, we first analyze a special case in Section III, where all subarrays share a uniform rotation angle, to gain insights into the impact of array rotation on mixed-field communication performance. Subsequently, in Section IV, we address the general case with subarray-wise rotation and propose an efficient double-layer algorithm to obtain a high-quality suboptimal solution to problem (P1).
@@
In this paper, we analytically investigate the impact of RAs on the performance of mixed-field communication systems, with a particular focus on the suppression of near-field inter-user interference as well as the unique mixed-field inter-user interference. We study the maximization of the sum-rate for near-field users in the presence of legacy far-field users. Furthermore, we extend our analysis to explore three additional system configurations:
  • Maximizing the sum-rate of far-field users while ensuring the communication rate requirements of each near-field user. For this configuration, the original optimization problem (P1) needs to incorporate an additional per-user rate constraint. The proposed algorithm remains directly applicable to solve this modified problem.
  • Maximizing the sum-rate of near-field users while ensuring that each far-field user meets their communication rate requirement. Similarly, this setup requires problem (P1) to incorporate per-user rate constraints for far-field users, and the proposed algorithm can be directly applied.
@@ III Analysis of Array Rotation on System Performance
In this section, we analytically characterize the effect of RAs on the rate performance of mixed-field communications. To gain useful insights, we begin by considering a special case where all subarrays share the same rotation angle, referred to as array rotation. For notational simplicity, we omit the subarray index $q$. Accordingly, we rewrite (II-B) as (III) at the bottom of the next page. From (III), it is clear that the system sum-rate is significantly affected by two types of interference: near-field inter-user interference, denoted by $|\mathbf{b}^{H}(\theta_{k},r_{k},\phi)\mathbf{b}(\theta_{i},r_{i},\phi)|^{2}$, and mixed-field interference from the far-field users to near-field users, denoted by $|\mathbf{b}^{H}(\theta_{k},r_{k},\phi)\mathbf{a}(\psi_{m},\phi)|^{2}$. Moreover, the analytical impact of array rotation on the system performance is intricately linked to these two types of interference. Therefore, we proceed by closely examining the influence of array rotation on each interference term in detail. To clearly characterize the effects of the rotation angle $\phi$ on these interference terms, we introduce the following definitions.
@@
The rotation-aware normalized near-field inter-user interference between any near-field users $k$ and $i$, $\forall{k,i\in\mathcal{K}}$ is defined as:
@@
The rotation-aware normalized mixed-field inter-user interference between the far-field user $m\in\mathcal{M}$ and near-field user $k\in\mathcal{K}$ is defined as:
@@
Moreover, it follows directly from (III) that the sum-rate is a monotonically decreasing function of both $\rho_{\rm NN}(\cdot)$ and $\rho_{\rm NF}(\cdot)$. Mathematically, these interference terms are explicitly given in (15) and (16), as shown at the top of the next page. These expressions, however, are complex and difficult to analyze directly. To address this, we obtain their closed-form expressions by using the Fresnel integrals, as presented below.
$\rho_{\rm NN}(\phi,\theta_{k},r_{k},\theta_{i},r_{i})=\frac{1}{N}\left|\sum_{n=-\tilde{N}}^{\tilde{N}}e^{j\frac{2\pi}{\lambda}\left[n^{2}\left(\frac{d^{2}\sin^{2}{(\phi-\theta_{k})}}{2r_{k}}-\frac{d^{2}\sin^{2}{(\phi-\theta_{i})}}{2r_{i}}\right)+n(d\cos{(\phi-\theta_{i})}-d\cos{(\phi-\theta_{k})})\right]}\right|.$ (15)
$\rho_{\rm NF}(\phi,\theta_{k},r_{k},\psi_{m})=\frac{1}{N}\left|\sum_{n=-\tilde{N}}^{\tilde{N}}e^{j\frac{2\pi}{\lambda}\left[n^{2}\frac{d^{2}\sin^{2}{(\phi-\theta_{k})}}{2r}+n(d\cos{(\psi_{m}-\phi)}-d\cos{(\phi-\theta_{k}))}\right]}\right|.$ (16)

### v4

@@ II-A2 Near-field channel model
For a typical near-field user, the channel is modeled based on spherical-wave propagation. Consider near-field user $k\in\mathcal{K}$, located at a distance of $r_{k}$ and at a spatial angle of $\theta_{k}$ from the center of the entire ULA. The coordinates of the $k$-th user are given by $\mathbf{r}_{k}=[r_{k}\cos{\theta_{k}},r_{k}\sin{\theta_{k}}]^{T}$. The distance from the $\bar{n}$-th antenna in subarray $q$ to user $k$, accounting for the subarray’s rotation, is:
$r^{(\bar{n})}_{k}=\|\mathbf{r}_{k}-\mathbf{w}_{q,\bar{n}}\|$
$=\sqrt{(r_{k}\cos{\theta_{k}}-c_{q}-\bar{n}d\cos{\phi_{q}})^{2}+(r_{k}\sin{\theta_{k}}-\bar{n}d\sin{\phi_{q}})^{2}}.$ (1)
Then, the near-field channel steering vector $\mathbf{b}_{q}(\theta_{k},r_{k},\phi_{q})\in\mathbb{C}^{\bar{N}}$ for subarray $q$ is given by
$\mathbf{b}_{q}(\theta_{k},r_{k},\phi_{q})=\left[e^{-j\frac{2\pi}{\lambda}(r^{(\bar{n})}_{k}-r_{k})}\right]_{n\in\mathcal{\bar{N}},q\in\mathcal{Q}}.\vskip-3.0pt$ (2)
@@
where $\bm{\phi}=[\phi_{-\tilde{Q}},\cdots,\phi_{\tilde{Q}}]^{T}$ denotes the vector of rotation angles of all subarrays. In this paper, we adopt a multipath channel model for the near-field users, comprising a line-of-sight (LoS) path and $L_{{\rm N},k}$ non-line-of-sight (NLoS) paths induced by scatterers in the environment. Then, the channel between the BS and near-field user $k$ is modeled as
$\mathbf{h}^{H}_{{\rm N},k}\!\!=\!\!\sqrt{N}{h}_{{\rm N},k}\mathbf{b}^{H}(\theta_{k},r_{k},\bm{\phi})\!+\!\sum^{L_{{\rm N},k}}_{\ell=1}\sqrt{N}{h}_{{\rm N},k,\ell}\mathbf{b}^{H}(\theta_{k,\ell},r_{k,\ell},\bm{\phi}),$ (4)
where ${h}_{{\rm N},k}$ and ${h}_{{\rm N},k,\ell}$ denote the complex-valued channel gains of the LoS path and the $\ell$-th NLoS path between the BS and near-field user $k$, respectively. [^3] For simplicity, we assume an omni-directional antenna gain model. Nonetheless, the proposed scheme can be readily extended to the directional antenna model described in [26, 29]. $\theta_{k,\ell}$ and $r_{k,\ell}$ are the angle and distance of the $\ell$-th scatterer associated with user $k$, measured relative to the center of the ULA.
@@ II-A3 Far-field channel model
The channel between the BS and far-field user $m\in\mathcal{M}$ follows the planar-wave propagation assumption. Thus, the far-field channel steering vector $\mathbf{a}_{q}(\psi_{m},\phi_{q})\in\mathbb{C}^{\bar{N}}$ for subarray $q$ is given by
$\mathbf{a}_{q}(\psi_{m},\phi_{q})=\left[e^{-j\frac{2\pi}{\lambda}\bar{n}d\cos{(\psi_{m}-\phi_{q})}}\right]_{\bar{n}\in\mathcal{\bar{N}},q\in\mathcal{Q}},\vskip-3.0pt$ (5)
where $\psi_{m}$ denotes the spatial angle of far-field user $m$. Accordingly, the overall far-field channel steering vector is given by
@@
Consequently, the multipath far-field channel between the BS and the far-field user $m\in\mathcal{M}$ can be expressed as
$\mathbf{h}^{H}_{{\rm F},m}=\sqrt{N}{h}_{{\rm F},m}\mathbf{a}^{H}(\psi_{m},\phi)+\sum^{L_{{\rm F},m}}_{\ell=1}\sqrt{N}{h}_{{\rm F},m,\ell}\mathbf{a}^{H}(\psi_{m,\ell},\phi),$ (7)
where ${h}_{{\rm F},m}$ and ${h}_{{\rm F},m,\ell}$ denote the complex-valued channel gains of the LoS path and the $\ell$-th NLoS path between the BS and user $m$, respectively.
@@ II-B Signal Model
Let $s_{{\rm N},k}\in\mathbb{C}$ and $s_{{\rm F},m}\in\mathbb{C}$ represent the signals transmitted by the BS to the $k$-th near-field user with power $P_{{\rm N},k}$ and to the $m$-th far-field user with power $P_{{\rm F},m}$, respectively. For legacy far-field users, the preconfigured beamforming vectors are defined as $\mathbf{w}_{{\rm F},m}\in\mathbb{C}^{N},m\in\mathcal{M}$. For the near-field users, to reduce power consumption and hardware cost in upper mid-band MIMO systems, a hybrid beamforming architecture is employed to serve the $K$ near-field users with $K$ $(K\ll N)$ radio frequency (RF) chains [12]. The transmitted signal vector is thus given by $\bar{\mathbf{s}}=\mathbf{F}_{\rm A}\mathbf{F}_{\rm D}\mathbf{s}$, where $\bar{\mathbf{s}}=[s_{{\rm N},1},\ldots,s_{{\rm N},K}]^{T}$, $\mathbf{F}_{\rm D}\in\mathbb{C}^{K\times K}$ is the digital precoder, and $\mathbf{F}_{\rm A}=[\mathbf{v}_{{\rm N},1},\ldots,\mathbf{v}_{{\rm N},K}]\in\mathbb{C}^{N\times K}$ is the analog precoder, with $\mathbf{v}_{{\rm N},k}\in\mathbb{C}^{N}$ representing the analog beamforming vector for the $k$-th near-field user. As such, the received signal at the $k$-th near-field user is given by [^4] To fully exploit the performance gains brought by RAs, we assume that perfect channel state information (CSI) for both near-field and far-field users is available at the BS [26]. In practice, their CSI can be effectively acquired through existing far-field and near-field beam training [31] and/or channel estimation techniques [32].
@@
where $\mathbf{w}_{{\rm N},i}=\mathbf{F}_{\rm A}\mathbf{f}_{{\rm D},i}$ with $\mathbf{f}_{{\rm D},i}$ being the $i$-th column of $\mathbf{F}_{\rm D}$. $n_{k}\sim\mathcal{CN}(0,\sigma^{2}_{k})$ represents the additive white Gaussian noise (AWGN) at the near-field user $k$. Accordingly, the achievable rate (bits/s/Hz) of near-field user $k$ is given by $R_{k}=\log_{2}\left(1+\gamma_{k}\right)$, where
@@ II-C Problem Formulation
To shed light onto the impact of RAs on mixed-field communication systems, we employ a low-complexity hybrid beamforming architecture [12]. In this architecture, $K$ beams are directed towards the near-field users using near-field beam-focusing, while multiple preconfigured maximum ratio transmission (MRT)-like beams are steered toward the legacy far-field users [11] with fixed power allocations. The beamforming vectors are defined as $\mathbf{v}_{{\rm N},k}=\frac{\mathbf{h}_{{\rm N},k}}{\|\mathbf{h}_{{\rm N},k}\|}$ for the $k$-th near-field user and $\mathbf{w}_{{\rm F},m}=\frac{\mathbf{h}_{{\rm F},m}}{\|\mathbf{h}_{{\rm F},m}\|}$ for the $m$-th far-field user. The digital beamforming is configured as an identity matrix, i.e., $\mathbf{F}_{\rm D}=\mathbf{I}_{K}$ [12, 33]. [^5] To enhance the system performance, the weighted minimum mean square error (WMMSE) or zero-forcing (ZF) based digital beamforming can be properly designed based on the analog precoder $\mathbf{F}_{\rm A}$ [12, 33], with performance evaluations presented in Section V. Consequently, the achievable rate of near-field user $k$ is given by (II-B) at the bottom of this page. We consider the system sum-rate maximization problem by jointly optimizing the power allocation for the near-field users $\{P_{{\rm N},k}\}$, and the subarray rotation angles $\bm{\phi}$. The optimization problem is formulated as follows:
$({\bf P1}):\penalty 10000\ \penalty 10000\ \penalty 10000\ \max_{\begin{subarray}{c}\{P_{{\rm N},k}\},\bm{\phi}\end{subarray}}$ $\penalty 10000\ \penalty 10000\ \sum_{k\in\mathcal{K}}\penalty 10000\ R_{k}\left(\{P_{{\rm N},k}\},\bm{\phi}\right)$ (11a)
s.t. $\penalty 10000\ \penalty 10000\ \sum_{k\in\mathcal{K}}P_{{\rm N},k}\leq P,$ (11b)
$\penalty 10000\ \penalty 10000\ \left|[\bm{\phi}]_{p}-[\bm{\phi}]_{q}\right|>\frac{\pi}{2},\penalty 10000\ \forall{p,q\in\mathcal{Q}},$ (11c)
$\penalty 10000\ \penalty 10000\ [\bm{\phi}]_{q}\in\mathcal{C}_{{\phi_{q}}},\penalty 10000\ \forall{q\in\mathcal{Q}},$ (11d)
@@
where (11b) enforces the BS transmit power constraint, with $P$ being the maximum transmit power. (11c) ensures that, for any two subarrays $p,q\in\mathcal{Q}$, their individual rotations $\bm{\phi}]_{p}$ and $\bm{\phi}]_{q}$ form an obtuse angle, thereby preventing mutual signal reflections. [21, 26]. (11d) restricts the rotation angle of subarray $q$ to its corresponding admissible range $\mathcal{C}_{\phi_{q}}=[\phi_{q,{\rm min}},\phi_{q,{\rm max}}]$.
@@
Optimization problem (P1) is non-convex due to the following factors: 1) The objective function is not jointly concave in terms of the optimization variables $\{P_{{\rm N},k}\}$ and $\bm{\phi}$; 2) these variables are intricately coupled within the objective function; and 3) the rotation angles $\bm{\phi}$ introduce complex dependencies in both the near-field and far-field steering vectors. These characteristics render the optimal solution of (P1) computationally challenging. To tackle this, we first analyze a special case in Section III, where all subarrays employ the same rotation angle, to gain analytical insight into the impact of array rotation on mixed-field communication performance. Subsequently, in Section IV, we address the general case with subarray-wise rotation and propose an efficient double-layer algorithm to obtain a high-quality suboptimal solution to problem (P1).
@@
In this paper, we analytically investigate the impact of RAs on the performance of mixed-field communication systems, with a particular focus on the suppression of near-field inter-user interference as well as the unique mixed-field inter-user interference. Specifically, we study the maximization of the sum-rate for near-field users in the presence of legacy far-field users. Furthermore, we extend our analysis to explore three additional system configurations:
  • Maximizing the sum-rate of the far-field users while ensuring the required communication rate for each near-field user. For this configuration, the original optimization problem (P1) needs to incorporate an additional per-user rate constraint. The proposed algorithm remains directly applicable to solve this modified problem.
  • Maximizing the sum-rate of the near-field users while ensuring that each far-field user meets their communication rate requirement. Similarly, this setup requires problem (P1) to incorporate per-user rate constraints for far-field users, and the proposed algorithm can be directly applied.
@@ III Analysis of Array Rotation on System Performance
In this section, we analytically characterize the effect of RAs on the rate performance of mixed-field communications. To gain useful insights, we begin by considering a special case where all subarrays share the same rotation angle, referred to as array rotation. For notational simplicity, we omit the subarray index $q$. Accordingly, we rewrite (II-B) as (III) at the bottom of this page. From (III), it is clear that the system sum-rate is significantly affected by two types of interference: near-field inter-user interference, represented by $|\mathbf{b}^{H}(\theta_{k},r_{k},\phi)\mathbf{b}(\theta_{i},r_{i},\phi)|^{2}$, and mixed-field interference from the far-field users to near-field users, represented by $|\mathbf{b}^{H}(\theta_{k},r_{k},\phi)\mathbf{a}(\psi_{m},\phi)|^{2}$. Moreover, the analytical impact of array rotation on system performance is intricately linked to these two types of interference. Therefore, we proceed by closely examining the influence of array rotation on each interference term in detail. To characterize the effects of the rotation angle $\phi\in[\phi_{\rm min},\phi_{\rm max}]$ on these interference terms, we introduce the following definitions.
@@
The rotation-dependent normalized near-field inter-user interference between any near-field users $k$ and $i$, $\forall{k,i\in\mathcal{K}}$ is defined as:
@@
The rotation-dependent normalized mixed-field inter-user interference between the far-field user $m\in\mathcal{M}$ and near-field user $k\in\mathcal{K}$ is defined as:
@@
Moreover, it follows directly from (III) that the sum-rate is a monotonically decreasing function of both $\rho_{\rm NN}(\cdot)$ and $\rho_{\rm NF}(\cdot)$. Mathematically, these interference terms are explicitly given in (15) and (16), as shown at the top of the next page. These expressions, however, are complex and difficult to analyze directly. To address this, we obtain closed-form expressions by using the Fresnel integrals, as presented below.
$\rho_{\rm NN}(\phi,\theta_{k},r_{k},\theta_{i},r_{i})=\frac{1}{N}\left|\sum_{n=-\tilde{N}}^{\tilde{N}}e^{j\frac{2\pi}{\lambda}\left[n^{2}\left(\frac{d^{2}\sin^{2}{(\phi-\theta_{k})}}{2r_{k}}-\frac{d^{2}\sin^{2}{(\phi-\theta_{i})}}{2r_{i}}\right)+n(d\cos{(\phi-\theta_{i})}-d\cos{(\phi-\theta_{k})})\right]}\right|.\vskip-3.0pt$ (15)
$\rho_{\rm NF}(\phi,\theta_{k},r_{k},\psi_{m})=\frac{1}{N}\left|\sum_{n=-\tilde{N}}^{\tilde{N}}e^{j\frac{2\pi}{\lambda}\left[n^{2}\frac{d^{2}\sin^{2}{(\phi-\theta_{k})}}{2r}+n(d\cos{(\psi_{m}-\phi)}-d\cos{(\phi-\theta_{k}))}\right]}\right|.\vskip-5.0pt$ (16)